OpenStudy (anonymous):

Write the equation 0.3x 2 + 5x - 7 = 0 in general form and then choose the value of "b." 70 50 5

OpenStudy (anonymous):

not sure what you mean we can multiply by 10 giving 3x^2 + 50x - 70 = 0 giving b = 50

OpenStudy (radar):

I hate problems where there is doubt, but I would multiply through by 10 getting: \[3x ^{2}+50x-70=0\] I believe that is the equation in general form, with "proper" coefficient for the x^2 term, if that is true then answer would be 50

OpenStudy (radar):

Great, joeywhite agrees.

OpenStudy (anonymous):

i've assumed that general form is ax^2 + bx + c = 0 where a,b and c are integers

OpenStudy (radar):

A reasonable assumption (I hope)

OpenStudy (anonymous):

lol - i agree

OpenStudy (radar):

Why they gave a option of 5, is just ornery. lol